how many gallons of a 50% antifreeze solution must be mixed with 60 gallons of 20% antifreeze to get a mixture that is 40% antifreeze
Answer by mananth(16946)

------ Percent ---------------- quantity
Antifreeze 20 ---------------- 60
Antifreeze 50 ---------------- x
Total 40 ---------------- 60 + x

20*60+50*x=40(60+x)
1200+50x=2400+40x
50x-40x=2400-1200
10x=1200
/10
x=120 gallons 50 % Antifreeze